# PLD: A Choice-Theoretic List-Wise Knowledge Distillation

### v2

@@ A.1 Definitions
Our goal is to compute the gradient s($\nabla_{s}\mathcal{L}(s)$, whose $i$-th component is $\frac{\partial\mathcal{L}}{\partial s_{i}}$. Due to the linearity of differentiation,
@@ 1. Derivative of the affine term:
The first term is $-\alpha_{k}s_{\pi^{*}_{k}}$.
@@ 2. Derivative of the log-sum-exp term:
A.3 Final Gradient s($\nabla_{s}\mathcal{L}(s)$:
@@
$\frac{\partial\mathcal{L}}{\partial s_{i}}$ $=\sum_{k=1}^{C}\frac{\partial L_{k}}{\partial s_{i}}$
$=\sum_{k=1}^{C}\alpha_{k}\left[\sigma_{k}(i)-\mathbf{1}\{i=\pi^{*}_{k}\}\right].$
@@ A.4 Convexity and Smoothness:
Each term $L_{k}(s)=\alpha_{k}\left[-s_{\pi^{*}_{k}}+\log\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})\right]$ is a sum of an affine function ($-\alpha_{k}s_{\pi^{*}_{k}}$) and a non-negatively weighted log-sum-exp function ($\alpha_{k}\phi_{k}(s)$). Affine functions are convex. The log-sum-exp function $\phi_{k}(s)$ is known to be convex. Assuming $\alpha_{k}\geq 0$ (which is true if they are derived from softmax probabilities, as in $q^{T}_{\pi^{*}_{k}}$), the term $\alpha_{k}\phi_{k}(s)$ is also convex. The sum of convex functions is convex, so each $L_{k}(s)$ is convex. The total loss $\mathcal{L}(s)=\sum_{k=1}^{C}L_{k}(s)$ is a sum of convex functions, and therefore, $\mathcal{L}(s)$ is convex in $s$.
@@ Appendix B Full Baseline Ablation
Table 6: Top-1 and Top-5 accuracy for baseline methods (ResNet-50 student, ResNet-152 teacher with Top-1 Acc. 79.61%). “-” indicates not applicable. CE: standard cross-entropy. LS: cross-entropy with label smoothing ($\epsilon=0.1$). KD: vanilla distillation with mixing weight $\alpha$ and temperature $\tau=2$. DIST: relational distillation with inter-class weight $\beta$, intra-class weight $\gamma$, and temperature $\tau=1$.
Method $\alpha$ $\beta$ $\gamma$ $\tau$ Top-1 Acc.Top-5 Acc.%
Oracle Teacher
Teacher (ResNet-152) - - - - 79.61 -
Baselines
CE - - - - 71.35 -
LS ($\epsilon=0.1$) - - - - 73.92 -
Vanilla KD ($\tau=2$)
KD 0.00 - - 2.00 75.92 92.82
0.10 - - 2.00 76.80 93.16
0.20 - - 2.00 76.65 93.11
0.30 - - 2.00 76.30 93.02
0.40 - - 2.00 76.33 93.10
0.50 - - 2.00 76.12 93.12
0.60 - - 2.00 75.94 92.79
0.70 - - 2.00 75.85 92.71
0.80 - - 2.00 75.21 92.39
0.90 - - 2.00 74.28 91.59
DIST ($\tau=1$)
DIST 1.00 2.00 2.00 1.00 75.77 92.67
@@ Appendix D Additional Experimental Setup
To comprehensively evaluate the relative merits of Plackett-Luce Distillation (PLD) compared to DIST and KD, we conduct experiments across multiple optimizers, various divergence weightings, and several logit-standardization schemes. We further quantify distributional alignment by measuring the KL divergence between student and teacher softmax outputs throughout training. Finally, to gain geometric insight into each loss, we visualize the loss landscapes of PLD, KD, and DIST.
@@ Appendix H Distribution Matching Analysis
Appendix I Loss Landscape Visualization
@@
I.1 Temperature Sensitivity of the PLD Loss Surface
@@
(a) $T=2.0$
(b) $T=1.0$
(c) $T=0.5$
(d) $T=0.1$

### v3

@@ A.1 Definitions
Our goal is to compute the gradient $\nabla_{s}\mathcal{L}(s)$, whose $i$-th component is $\frac{\partial\mathcal{L}}{\partial s_{i}}$. Due to the linearity of differentiation,
@@ 1. Derivative of the affine term:
The first term is $\alpha_{k}s_{\pi^{*}_{k}}$.
@@ 2. Derivative of the log-sum-exp term:
A.3 Final Gradient $\nabla_{s}\mathcal{L}(s)$:
@@
$\frac{\partial\mathcal{L}}{\partial s_{i}}$ $=\sum_{k=1}^{C}\frac{\partial L_{k}}{\partial s_{i}}$ $=\sum_{k=1}^{C}\alpha_{k}\left[\sigma_{k}(i)-\mathbf{1}\{i=\pi^{*}_{k}\}\right].$
@@ A.4 Convexity and Smoothness:
Each term $L_{k}(s)=\alpha_{k}\left[-s_{\pi^{*}_{k}}+\log\sum_{\ell=k}^{C}\exp(s_{\pi^{*}_{\ell}})\right]$ is a sum of an affine function ($\alpha_{k}s_{\pi^{*}_{k}}$) and a non-negatively weighted log-sum-exp function ($\alpha_{k}\phi_{k}(s)$). Affine functions are convex. The log-sum-exp function $\phi_{k}(s)$ is known to be convex. Assuming $\alpha_{k}\geq 0$ (which is true if they are derived from softmax probabilities, as in $q^{T}_{\pi^{*}_{k}}$), the term $\alpha_{k}\phi_{k}(s)$ is also convex. The sum of convex functions is convex, so each $L_{k}(s)$ is convex. The total loss $\mathcal{L}(s)=\sum_{k=1}^{C}L_{k}(s)$ is a sum of convex functions, and therefore, $\mathcal{L}(s)$ is convex in $s$.
@@ Appendix B Full Baseline Ablation
Table 8: Hyperparameter grid search on ImageNet-1K. Each configuration uses a ResNet-50 student distilled from a ResNet-152 teacher (Top-1 Acc. 79.61%). "–" indicates not applicable. The best setting in each group is highlighted.
Method Hyperparameters Accuracy (%)
$\alpha$ $\beta$ $\gamma$ $\tau$ Top-1 Top-5
Teacher 79.61
CE 71.35
LS ($\epsilon{=}0.1$) 73.92
KD hyperparameter sweep ($\tau{=}2$)
0.00 2.00 75.92 92.82
KD 0.10 2.00 76.80 93.16
0.20 2.00 76.65 93.11
0.30 2.00 76.30 93.02
0.40 2.00 76.33 93.10
0.50 2.00 76.12 93.12
0.60 2.00 75.94 92.79
0.70 2.00 75.85 92.71
0.80 2.00 75.21 92.39
0.90 2.00 74.28 91.59
DIST hyperparameter sweep ($\tau{=}1$)
1.00 2.00 2.00 1.00 75.77 92.67
@@ Appendix D Additional Experimental Setup
To comprehensively evaluate the relative merits of Plackett-Luce Distillation (PLD) compared to DIST and KD, we conduct experiments across multiple optimizers, various divergence weightings, runtime and several logit-standardization schemes. We further quantify distributional alignment by measuring the KL divergence between student and teacher softmax outputs throughout training. Finally, to gain geometric insight into each loss, we visualize the loss landscapes of PLD, KD, and DIST.
@@ Appendix H Distribution Matching Analysis
Table 13: Total training time (minutes) for all CIFAR-100 experiments. PLD achieves comparable efficiency to KD, DIST, and DKD despite its list-wise formulation.
Model (Teacher Student) CE PLD KD DIST DKD AT FitNet PKT RKD SP VID
[…]
ResNet50 MobileNetV2 23.62 102.59 102.23 102.49 102.54 108.17 113.18 102.28 128.65 102.78 132.06
ResNet56 ResNet20 22.97 33.42 33.57 33.81 33.46 34.55 33.30 33.89 36.21 33.67 36.57
WideResNet-40-2 WideResNet-40-1 27.67 33.10 37.56 36.73 34.27 36.77 38.05 36.84 37.62 37.01 37.35
Appendix I Runtime Analysis
We evaluate the computational efficiency of PLD compared with other distillation methods discussed in Section 5. PLD involves a single sorting operation to obtain the teacher-optimal permutation, similar to ListMLE xia2008listwise ; therefore, its complexity is only $O(C\log C)$ far lower than the $O(C!)$ enumeration of all possible rankings. In practice, this adds negligible overhead: PLD trains at nearly the same speed as logit-based methods such as KD, DIST, and DKD. Table 13 reports the total training time (in minutes) for all CIFAR-100 experiments in Table 1. Despite its list-wise formulation, PLD’s runtime remains comparable to KD and DIST, and substantially faster than other methods such as RKD or VID.
Appendix J Loss Landscape Visualization
(a) $T=2.0$
(b) $T=1.0$
(c) $T=0.5$
(d) $T=0.1$
@@ Appendix I Runtime Analysis
J.1 Temperature Sensitivity of the PLD Loss Surface
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. 2.
    Limitations
  5. Question: Does the paper discuss the limitations of the work performed by the authors?
  6. Answer: [Yes]
  7. 3.
    Theory assumptions and proofs
  8. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  9. Answer: [Yes]
  10. 4.
    Experimental result reproducibility
  11. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  12. Answer: [Yes]
  13. 5.
    Open access to data and code
  14. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  15. Answer: [Yes]
  16. 6.
    Experimental setting/details
  17. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  18. Answer: [Yes]
  19. 7.
    Experiment statistical significance
  20. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  21. Answer: [Yes]
  22. Justification: We report mean Top-1 accuracies on CIFAR-100 along with their corresponding standard errors computed over three independent runs with different random seeds.
  23. 8.
    Experiments compute resources
  24. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  25. Answer: [Yes]
  26. 9.
    Code of ethics
  27. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  28. Answer: [Yes]
  29. 10.
    Broader impacts
  30. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  31. Answer: [N/A]
  32. Justification: This work is foundational on model compression and does not present a direct application or risk. Indirectly, more efficient models may reduce energy and hardware costs, but we do not see specific societal harms arising from the method itself.
  33. 11.
    Safeguards
  34. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  35. Answer: [N/A]
  36. Justification: No high-risk or dual-use data/models are released in this work
  37. 12.
    Licenses for existing assets
  38. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  39. Answer: [Yes]
  40. 13.
    New assets
  41. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  42. Answer: [N/A]
  43. Justification: We do not release any new datasets or large models with this submission.
  44. 14.
    Crowdsourcing and research with human subjects
  45. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  46. Answer: [N/A]
  47. Justification: No human-subject data or crowdsourcing is involved.
  48. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  49. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  50. Answer: [N/A]
  51. Justification: This work does not involve human subjects.
  52. 16.
    Declaration of LLM usage
  53. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  54. Answer: [N/A]
  55. Justification: No large language models are used in our core methodology.